Suppose the demand and supply curves for a product is given by


𝑄 = 500 βˆ’ 2𝑃


𝑄 = βˆ’100 + 3𝑃


a) Which is the supply curve and why?


b) Graph the demand and supply curves.


c) Compute the equilibrium price and quantity


d) If the current price is 100, what is the quantity demanded and quantity supplied? How would you describe this situation and what would you expect to happen in this market?


e) Suppose that the demand changes to 𝑄 = 600 βˆ’ 2𝑃. Find the new equilibrium price and quantity and show this on your graph.

1
Expert's answer
2022-03-07T12:09:37-0500

(a) "Q=-100+3p" is the supply curve since it is an increasing function and supply curve is also increasing in nature. The more you put value of p the more will be Q.


(b)



(c) at equilibrium, Supply = Demand

"Qs = Qd"

so, "500 \u2212 2\ud835\udc43 = \u2212100 + 3\ud835\udc43"

hence equilibrium price is 120 and equilibrium quantity is 260


(d) If price is 100, then , "Qd = 500 - 2(100) = 300" and "Qs = -100 + 3(100) = 200" .

This is described a case of excess demand. The effect is that it will cause an upward pressure in prices.


(e) at equilibrium, "Qs = Qd"

so, "600 \u2212 2\ud835\udc43 = \u2212100 + 3\ud835\udc43"

hence equilibrium price is 140 and equilibrium quantity is 320. This would cause a shift in the demand curve to the right as shown in the graph in (b) above.